# Topological orbital Hall effect caused by skyrmions and antiferromagnetic skyrmions

### v1

@@ Introduction
Besides this true antiferromagnetic skyrmion [21, 23, 24], synthetic antiferromagnetic skyrmions (cf. Fig. 4b) have been considered as well [22, 26, 25]. In the latter case, the interpretation of two coexisting sub-skyrmions is even more applicable. This is especially true once we extend the sd-Hamiltonian to first and second nearest neighbor hopping and consider only second-nearest neighbor hopping $t_{2}$.
The corresponding charge, orbital and spin Hall conductivites for $t_{1}=0$ and $t_{2}=-1\,\mathrm{eV}$ are shown in Fig. 4c-e. The bands are doubly degenerate resembling the opposite spin alignment in the two sublattices hosting skyrmions with opposite emergent magnetic fields. As a consequence, the topological Hall effect is compensated for every energy (Fig. 4c) but the orbital and spin Hall conductivities are finite (Fig. 4d,e). Their energy dependencies are similar to those of the conventional skyrmion presented in Fig. 2c,d but their magnitude is larger due to the two coexisting sublattices and the larger skyrmion size. Qualitatively similar results can be obtained for nearest-neighbor hopping only ($t_{1}=-1\,\mathrm{eV}$ and $t_{2}=0$, presented in Fig. S2d-f of the Supplementary Material) and both hoppings considered ($t_{1}=-2/3\,\mathrm{eV}$ and $t_{2}=-2/3\,\mathrm{eV}$, presented in Fig. S2a-c of the Supplementary Material).
@@ Conclusions
In summary, we have shown that skyrmions and antiferromagnetic skyrmions give rise to a topological orbital Hall effect. For conventional skyrmions, the net emergent field forces electrons onto circular trajectories and causes skipping orbits at the edges akin to the quantum Hall effect. This gives rise to a topological Hall effect for which the charge currents are spin and orbital polarized. An antiferromagnetic skyrmions on the other hand, consists of two oppositely oriented subskyrmions on two sublattices. The opposite emergent fields cause opposite topological Hall effects and opposite spin and orbital polarization giving rise to a compensated topological Hall effect but net spin and orbital Hall effects.
Since the orbital angular momentum can take arbitrarily high values, the orbital Hall conductivity is much larger than the spin Hall conductivity. The orbital Hall conductivity scales roughly quadratically with the skyrmion diameter while the spin and charge Hall conductivities scale roughly linearly. This means these textures could serve as generators of large orbital currents that can potentially transport information and give rise to considerable orbital torques [46, 47].
@@ Methods
A skyrmion centered at $\boldsymbol{r}_{0}=0$ is modelled by
$\boldsymbol{m}=\begin{pmatrix}x\sin(2\pi r/\lambda)/r\\ y\sin(2\pi r/\lambda)/r\\ \cos(2\pi r/\lambda)\end{pmatrix}$ (2)
for $r=\sqrt{x^{2}+y^{2}}<\lambda/2$ and $\boldsymbol{m}=-\boldsymbol{e}_{z}$ otherwise.
@@
$N_{\mathrm{Sk}}=\frac{1}{4\pi}\int_{xy}\boldsymbol{m}(\boldsymbol{r})\cdot\left[\frac{\partial\boldsymbol{m}(\boldsymbol{r})}{\partial x}\times\frac{\partial\boldsymbol{m}(\boldsymbol{r})}{\partial y}\right]\,\mathrm{d}^{2}r.$ (3)
@@
$S_{z,\nu}(\boldsymbol{k})=\frac{\hbar}{2}\langle\nu\boldsymbol{k}|\sigma_{z}|\nu\boldsymbol{k}\rangle.$ (4)
@@
$\langle\nu\boldsymbol{k}|L_{z}|\alpha\boldsymbol{k}\rangle=i\frac{e\hbar^{2}}{4g_{L}\mu_{\mathrm{B}}}\sum_{\beta\neq\nu,\alpha}\left(\frac{1}{E_{\beta\boldsymbol{k}}-E_{\nu\boldsymbol{k}}}+\frac{1}{E_{\beta\boldsymbol{k}}-E_{\alpha\boldsymbol{k}}}\right)$
×$\times\left(\langle\nu\boldsymbol{k}|v_{x}|\beta\boldsymbol{k}\rangle\langle\beta\boldsymbol{k}|v_{y}|\alpha\boldsymbol{k}\rangle-\langle\nu\boldsymbol{k}|v_{y}|\beta\boldsymbol{k}\rangle\langle\beta\boldsymbol{k}|v_{x}|\alpha\boldsymbol{k}\rangle\right).$ (5)
Here, $\boldsymbol{v}=\frac{1}{\hbar}\nabla_{\boldsymbol{k}}H$ is the velocity operator. $L_{\nu,z}(\boldsymbol{k})$ are the diagonal elements of the tensor $L_{\nu,z}(\boldsymbol{k})=\langle\nu\boldsymbol{k}|L_{z}|\nu\boldsymbol{k}\rangle$.
@@
$\sigma_{xy}(E_{\text{F}})=-\frac{e^{2}}{h}\sum_{\nu}\frac{1}{2\pi}\int_{E_{\nu\boldsymbol{k}}\leq E_{\text{F}}}\Omega_{\nu,z}(\boldsymbol{k})\,\mathrm{d}^{2}k$ (6)
@@
$\Omega_{\nu,z}(\boldsymbol{k})=-2\hbar^{2}\ \text{Im}\ \sum_{\mu\neq\nu}\frac{\langle\nu\boldsymbol{k}|v_{x}|\mu\boldsymbol{k}\rangle\langle\mu\boldsymbol{k}|v_{y}|\nu\boldsymbol{k}\rangle}{(E_{\nu\boldsymbol{k}}-E_{\mu\boldsymbol{k}})^{2}}.$ (7)
The orbital and spin Hall conductivities [57]
$\sigma^{L_{z}}_{xy}(E_{\text{F}})$ $=\frac{e}{\hbar}\sum_{\nu}\frac{1}{(2\pi)^{2}}\int_{E_{\nu\boldsymbol{k}}\leq E_{\text{F}}}\Omega_{\nu,z}^{L_{z}}(\boldsymbol{k})\,\mathrm{d}^{2}k,$ (8)
$\sigma^{S_{z}}_{xy}(E_{\text{F}})$ $=\frac{e}{\hbar}\sum_{\nu}\frac{1}{(2\pi)^{2}}\int_{E_{\nu\boldsymbol{k}}\leq E_{\text{F}}}\Omega_{\nu,z}^{L_{z}}(\boldsymbol{k})\,\mathrm{d}^{2}k$ (9)
@@
$\Omega_{\nu,z}^{L_{z}}(\boldsymbol{k})$ $=-2\hbar^{2}\ \text{Im}\ \sum_{\mu\neq\nu}\frac{\langle\nu\boldsymbol{k}|j_{x}^{L_{z}}|\mu\boldsymbol{k}\rangle\langle\mu\boldsymbol{k}|v_{y}|\nu\boldsymbol{k}\rangle}{(E_{\nu\boldsymbol{k}}-E_{\mu\boldsymbol{k}})^{2}},$ (10)
$\Omega_{\nu,z}^{S_{z}}(\boldsymbol{k})$ $=-2\hbar^{2}\ \text{Im}\ \sum_{\mu\neq\nu}\frac{\langle\nu\boldsymbol{k}|j_{x}^{S_{z}}|\mu\boldsymbol{k}\rangle\langle\mu\boldsymbol{k}|v_{y}|\nu\boldsymbol{k}\rangle}{(E_{\nu\boldsymbol{k}}-E_{\mu\boldsymbol{k}})^{2}},$ (11)
where $\langle\nu\boldsymbol{k}|j_{x}^{L_{z}}|\mu\boldsymbol{k}\rangle=\frac{1}{2}\sum_{\alpha}[\langle\nu\boldsymbol{k}|v_{x}|\alpha\boldsymbol{k}\rangle\langle\alpha\boldsymbol{k}|L_{z}|\mu\boldsymbol{k}\rangle+\langle\nu\boldsymbol{k}|L_{z}|\alpha\boldsymbol{k}\rangle\langle\alpha\boldsymbol{k}|v_{x}|\mu\boldsymbol{k}\rangle]$ is the orbital current operator and $\langle\nu\boldsymbol{k}|j_{x}^{S_{z}}|\mu\boldsymbol{k}\rangle=\frac{1}{2}[\langle\nu\boldsymbol{k}|v_{x}|\nu\boldsymbol{k}\rangle\langle\mu\boldsymbol{k}|S_{z}|\mu\boldsymbol{k}\rangle+\langle\nu\boldsymbol{k}|S_{z}|\nu\boldsymbol{k}\rangle\langle\mu\boldsymbol{k}|v_{x}|\mu\boldsymbol{k}\rangle]$ the spin current operator.
@@
Data that support the findings of this work are available from the authors on reasonable request.
@@ Acknowledgements
B.G. and L.S performed the calculations. B.G. wrote the manuscript with significant inputs from all authors. B.G. prepared the figures. All authors discussed the results. B.G. planned the project. B.G. and I.M. supervised the project. I.M. provided the funding.

### v2

@@ Introduction
Besides this true antiferromagnetic skyrmion [21, 23, 24], synthetic antiferromagnetic skyrmions (cf. Fig. 5b) have been considered as well [22] and have even been observed experimentally in multilayer systems [26, 25]. In the latter case, the interpretation of two coexisting sub-skyrmions is even more applicable. This is especially true once we extend the $s$-$d$ Hamiltonian to first-nearest neighbor hopping with amplitude $t_{1}$ (the hopping between the different sublattices) and second-nearest neighbor hopping with amplitude $t_{2}$ (the hopping within the sublattices) and consider only second-nearest neighbor hopping $t_{2}$.
The corresponding charge, orbital and spin Hall conductivites for $t_{1}=0$ and $t_{2}=-1\,\mathrm{eV}$ are shown in Fig. 5c-e. The bands are doubly degenerate resembling the opposite spin alignment in the two sublattices hosting skyrmions with opposite emergent magnetic fields. As a consequence, the topological Hall effect is compensated for every energy (Fig. 5c) but the orbital and spin Hall conductivities are finite (Fig. 5d,e). Their energy dependencies are similar to those of the conventional skyrmion presented in Fig. 2c,d but their magnitude is larger due to the two coexisting sublattices and the larger skyrmion size. Qualitatively similar results can be obtained for nearest-neighbor hopping only ($t_{1}=-1\,\mathrm{eV}$ and $t_{2}=0$, presented in Fig. S5a-c of the Supplementary Material) and both hoppings considered ($t_{1}=-2/3\,\mathrm{eV}$ and $t_{2}=-2/3\,\mathrm{eV}$, presented in Fig. S5d-f of the Supplementary Material).
@@
Fig. S6 in the Supplementary Material shows equivalent results for an antiferromagnetic bimeron [92, 28, 96]. Like for the antiferromagnetic skyrmion discussed above, the topological Hall effect vanishes due to a compensated topological charge. However, since the two bimerons do not have a net out-of-plane moment, the spin Hall effect is zero as well. Therefore, the antiferromagnetic bimeron, that has been observed experimentally [97, 98], might be the ideal platform to investigate the topological orbital Hall effect, as it exhibits a pure orbital Hall effect that is not superimposed by a charge or spin Hall effect.
@@ Conclusions
In summary, we have shown that skyrmions and antiferromagnetic skyrmions give rise to a topological orbital Hall effect. For conventional skyrmions, the net emergent field forces electrons onto circular trajectories and causes skipping orbits at the edges akin to the quantum Hall effect. This gives rise to a topological Hall effect for which the charge currents are spin and orbital polarized. An antiferromagnetic skyrmions on the other hand, consists of two oppositely oriented subskyrmions on two sublattices. The opposite emergent fields cause opposite topological Hall effects and opposite spin and orbital polarization giving rise to a compensated topological Hall effect but net spin and orbital Hall effects. An antiferromagnetic bimeron even exhibits a pure orbital Hall effect.
Since the orbital angular momentum is not restricted by a quantum number but increases with the skyrmion area, the orbital Hall conductivity is much larger than the spin Hall conductivity. The orbital Hall conductivity scales roughly quadratically with the skyrmion area while the spin and charge Hall conductivities scale roughly linearly (cf. Fig. S2 of the Supplementary Material). This means these textures could serve as generators of large orbital currents that can potentially transport information and give rise to considerable orbital torques [46, 47].
@@
It is worth noting again that in the present study we have focused on the intrinsic contribution of a skyrmionic texture to the orbital Hall effect in an easy-to-understand model system. We have restricted the model to $s$ orbitals, have disregarded spin-orbit coupling and did not consider scattering at defects. Taking these effects into account in a future study could be interesting as additional effect like the anomalous Hall effect or extrinsic effects like the ‘side jump’ and ‘skew scattering’ might occur that can be important corrections [75, 77]. Furthermore, it has been shown recently that the anomalous position has to be considered in the definition of the velocity operator once the system includes more than just $s$ orbitals [73]. As explained in Ref. [76], a complete quantum mechanical theory of non-equilibrium orbital angular momentum dynamics is not yet available and has to be derived from the ground up in the future.
@@ Methods
A skyrmion centered at $\mathbf{r}_{0}=0$ is modelled by
$\mathbf{m}=\begin{pmatrix}x\sin(2\pi r/\lambda)/r\\ y\sin(2\pi r/\lambda)/r\\ \cos(2\pi r/\lambda)\end{pmatrix}$ (2)
for $r=\sqrt{x^{2}+y^{2}}<\lambda/2$ and $\mathbf{m}=-\mathbf{e}_{z}$ otherwise.
@@
$N_{\mathrm{Sk}}=\frac{1}{4\pi}\int_{xy}\mathbf{m}(\mathbf{r})\cdot\left[\frac{\partial\mathbf{m}(\mathbf{r})}{\partial x}\times\frac{\partial\mathbf{m}(\mathbf{r})}{\partial y}\right]\,\mathrm{d}^{2}r.$ (3)
@@
$S_{z,\nu}(\mathbf{k})=\frac{\hbar}{2}\langle\nu\mathbf{k}|\sigma_{z}|\nu\mathbf{k}\rangle.$ (4)
@@
$\langle\nu\mathbf{k}|L_{z}|\alpha\mathbf{k}\rangle=\mathrm{i}\frac{e\hbar^{2}}{4g_{L}\mu_{\mathrm{B}}}\sum_{\beta\neq\nu,\alpha}\left(\frac{1}{E_{\beta\mathbf{k}}-E_{\nu\mathbf{k}}}+\frac{1}{E_{\beta\mathbf{k}}-E_{\alpha\mathbf{k}}}\right)$
×$\times\left(\langle\nu\mathbf{k}|v_{x}|\beta\mathbf{k}\rangle\langle\beta\mathbf{k}|v_{y}|\alpha\mathbf{k}\rangle-\langle\nu\mathbf{k}|v_{y}|\beta\mathbf{k}\rangle\langle\beta\mathbf{k}|v_{x}|\alpha\mathbf{k}\rangle\right).$ (5)
Here, $\mathbf{v}=\frac{1}{\hbar}\nabla_{\mathbf{k}}H$ is the velocity operator. $L_{\nu,z}(\mathbf{k})$ are the diagonal elements of the tensor $L_{\nu,z}(\mathbf{k})=\langle\nu\mathbf{k}|L_{z}|\nu\mathbf{k}\rangle$. Note that we have corrected a mistake in Ref. [57] where $\mathrm{Im}$ was used instead of the imaginary unit $i$.
@@
$\sigma_{xy}(E_{\text{F}})=-\frac{e^{2}}{h}\sum_{\nu}\frac{1}{2\pi}\int_{E_{\nu\mathbf{k}}\leq E_{\text{F}}}\Omega_{\nu,z}(\mathbf{k})\,\mathrm{d}^{2}k$ (6)
@@
$\Omega_{\nu,z}(\mathbf{k})=-2\hbar^{2}\ \text{Im}\ \sum_{\mu\neq\nu}\frac{\langle\nu\mathbf{k}|v_{x}|\mu\mathbf{k}\rangle\langle\mu\mathbf{k}|v_{y}|\nu\mathbf{k}\rangle}{(E_{\nu\mathbf{k}}-E_{\mu\mathbf{k}})^{2}}.$ (7)
The intrinsic orbital and spin Hall conductivities [57]
$\sigma^{L_{z}}_{xy}(E_{\text{F}})$ $=\frac{e}{\hbar}\sum_{\nu}\frac{1}{(2\pi)^{2}}\int_{E_{\nu\mathbf{k}}\leq E_{\text{F}}}\Omega_{\nu,z}^{L_{z}}(\mathbf{k})\,\mathrm{d}^{2}k,$ (8)
$\sigma^{S_{z}}_{xy}(E_{\text{F}})$ $=\frac{e}{\hbar}\sum_{\nu}\frac{1}{(2\pi)^{2}}\int_{E_{\nu\mathbf{k}}\leq E_{\text{F}}}\Omega_{\nu,z}^{S_{z}}(\mathbf{k})\,\mathrm{d}^{2}k$ (9)
@@
$\Omega_{\nu,z}^{L_{z}}(\mathbf{k})$ $=-2\hbar^{2}\ \text{Im}\ \sum_{\mu\neq\nu}\frac{\langle\nu\mathbf{k}|j_{x}^{L_{z}}|\mu\mathbf{k}\rangle\langle\mu\mathbf{k}|v_{y}|\nu\mathbf{k}\rangle}{(E_{\nu\mathbf{k}}-E_{\mu\mathbf{k}})^{2}},$ (10)
$\Omega_{\nu,z}^{S_{z}}(\mathbf{k})$ $=-2\hbar^{2}\ \text{Im}\ \sum_{\mu\neq\nu}\frac{\langle\nu\mathbf{k}|j_{x}^{S_{z}}|\mu\mathbf{k}\rangle\langle\mu\mathbf{k}|v_{y}|\nu\mathbf{k}\rangle}{(E_{\nu\mathbf{k}}-E_{\mu\mathbf{k}})^{2}},$ (11)
where $\langle\nu\mathbf{k}|j_{x}^{L_{z}}|\mu\mathbf{k}\rangle=\frac{1}{2}\sum_{\alpha}[\langle\nu\mathbf{k}|v_{x}|\alpha\mathbf{k}\rangle\langle\alpha\mathbf{k}|L_{z}|\mu\mathbf{k}\rangle+\langle\nu\mathbf{k}|L_{z}|\alpha\mathbf{k}\rangle\langle\alpha\mathbf{k}|v_{x}|\mu\mathbf{k}\rangle]$ is the orbital current operator and $\langle\nu\mathbf{k}|j_{x}^{S_{z}}|\mu\mathbf{k}\rangle=\frac{1}{2}[\langle\nu\mathbf{k}|v_{x}|\nu\mathbf{k}\rangle\langle\mu\mathbf{k}|S_{z}|\mu\mathbf{k}\rangle+\langle\nu\mathbf{k}|S_{z}|\nu\mathbf{k}\rangle\langle\mu\mathbf{k}|v_{x}|\mu\mathbf{k}\rangle]$ the spin current operator.
@@
The data that support the findings of this work are available at DOI: 10.5281/zenodo.13919926.
@@ Acknowledgements
B.G. and L.S. performed calculations. B.G. wrote the manuscript with significant inputs from all authors. B.G. prepared the figures. All authors discussed the results. B.G. planned the project. B.G. and I.M. supervised the project. I.M. provided the funding.